[figure: 9.7 cm]

MAGNON AND SOLITON EXCITATIONS IN A CARRIER-POOR, ONE-DIMENSIONAL S = 1/2 ANTIFERROMAGNET

F. Steglich, M. Lang, P. Thalmeier, B. Schmidt, M. Köppen, R. Helfrich,

1B. Wand, ^{2}A. Ochiai Max-Planck Institute for Chemical Physics of Solids, 01 187 Dresden, Germany

1Institute for Solid-State Physics, Univ. of Technology, 64289 Darmstadt, Germany 2Department of Material Science and Technology, Niigita Univ., Niigita 950-21 , Japan

Upon cooling to below T_{C} = 293K, the compound Yb_{4}As_{3} undergoes a structural transition from a cubic high-temperature phase into a trigonal low-T structure. This transition is accompanied by a change in the electronic structure from a homogeneous intermediatevalence metallic state, with an Yb valence close to 2.25, to a low-T semimetallic phase, with a carrier concentration n ≈ 10^{−3} / formula unit. The low-T phase can be characterized by a charge ordering of Yb^{3+} ions along the < 111 > direction. As a result of a superexchange interaction among the Yb^{3+} spins mediated by the surrounding As atoms, S = 1/2 antiferromagnetic (afm) Heisenberg chains are formed that are coupled only very weakly to each other. Here we report measurements of the specific heat, thermal expansion and thermal conductivity on Yb_{4}As_{3} aiming at an investigation of its low-temperature magnetic properties. A large in-T linear specific heat, C = γT, γ = 200mJ/K^{2}mole, is observed for B = 0 and T < 2K. This term is completely independent of the carrier concentration, which can be strongly modified by partial substitution of either P or Sb for As. We attribute γ to the magnon excitations of the afm chains. In finite magnetic fields, distinct anomalies were found in the above thermodynamic and transport quantities. These anomalies are well described by the classical sine-Gordon soliton solutions for an easy–plane Heisenberg antiferromagnet. Our findings strongly suggest that Yb_{4}As_{3} represents the first example of an afm S = 1/2 spin–chain system where this type of non–linear excitations could be identified.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : F. Steglich

Address for correspondence : Max-Planck Institute for Chemical Physics of Solids, 01 187 Dresden, Germany

Email address : steglich@cpfs.mpg.de